# Generate E-commerce Product Background by Integrating Category Commonality and Personalized Style

### v1

@@ 4.2 Dataset Construction and Analysis
Figure 4: Examples in BG60k. Compared with images in Fig. 3, the training images in BG60k are more appealing and exquisite.
The image number of the categories in BG60K follows the long-tail challenging distribution of the real advertising images with high impressions, since there are around $6\%$ of categories containing over 100 images, while over $50\%$ of categories less than 10 images. The detailed distribution of categories in BG60k could be found in the supplementary materials. We also visualize some examples of BG60K in Fig. 4, which proves that the training images are exquisite and exhibit prominent products.
5 Experiments
5.1 Implementation Details
Our model is trained on BG60k for 300 epochs with a learning rate 1e-5. The total batch size is 128 and each image is resized to $512\times 512$. The identifier is implemented with the combination of sks [18] and the id of the category. For data augmentation, the size of $\theta$ is 50, and the $\sigma$ in mixup is sampled from a uniform distribution $\mathcal{U}(0.75,0.95)$. Other hyper-parameters of the training are consistent with ControlNet. During inference, we adopt a 50-step DDIM sampler with the classifier-free guidance scale 2.0.
5.2 Evaluation Metrics
Quantitative Metrics: The generated backgrounds are evaluated from two aspects. On the one hand, the background should be exquisite, either photorealistic (for products like home appliances) or artistically stylized (for products like cosmetics). It should also be harmonious with the product to make the entire image look like an advertising image. Following Yang et al. [26], we measure this aspect by FID score [6], which reflects the distribution gap between features of two groups of images. On the other hand, the background of the generated image is supposed to be similar to the reference image (if given). This is quantified by the similarity of CLIP visual features of the two images.
User Study: We further conduct the user study for a more comprehensive comparison. Concretely, we sample one hundred product-reference pairs from BG-pair, and combine the original product, the reference image, and images generated by our method and counterparts. The order of the generated images is shuffled randomly for each group. Then we invite twenty experienced advertising practitioners to rank the consistency with the reference image and the quality of the generated background. When evaluating the background consistency, the practitioners are mainly concerned about the layout, elements, color and overall style. When evaluating the image quality, the practitioners are particularly concerned about the overall aesthetics, the harmony and the suitability of serving as the advertising image.
5.3 Comparison with State-of-the-Art
5.3.1 Large-Scale Background Generation
In this subsection, we evaluate our method for efficient large-scale background generation. Without extra condition input (the reference image) included, we remove PG and compare CG with the state-of-the-art inpainting methods Stable Diffusion and ControlNet, as well as an unconditional inpainting method LaMa [20] on BG1k. Since it is not feasible to carefully devise the prompt for each product, the backgrounds are generated based on the prompt template. Specifically, we adopt ’A photo of [category], in the background of [D]’ as the prompt template for CG, and ’A photo of [category]’ for Stable Diffusion and ControlNet, as [D] is the garbled character for them.
@@ 5.3.1 Large-Scale Background Generation
Table 1 summarizes the comparison on CLIP similarity (the generated image v.s. the original image) and FID. It could be found that CG has the superiority of 1.25 CLIP similarity and 1.85 FID. It means that the backgrounds generated by CG are more like the original high-quality backgrounds and thus better meet the category-wise requirements of the advertisements in the large-scale generation.
@@
Table 1: Comparison with image inpainting methods conditioned on the prompt template on BG1k.
Figure 5: (a) tSNE visualization, where the circle represents the cluster center in training data. Each triangle/square indicates the background embedding of one image generated by CG/ControlNet. (b) Visualization of comparison on the category refrigerator. Compared with ControlNet, CG could generate more realistic and exquisite backgrounds.
For a more comprehensive analysis on whether our generated backgrounds contain the category commonality, we visualize the features of backgrounds from our method and ControlNet by tSNE in Fig. 5 (a). It could be seen that the features of backgrounds generated by CG are clustered around the corresponding centers more compactly. On the contrary, background features generated by ControlNet are farther away from the centers, and even contain some outliers. Furthermore taking the category ’refrigerator’ as an example, we compare the generated images in Fig. 5 (b). It turns out that ControlNet might generate backgrounds lacking photorealism (the first column), irrelevant to the product (the third column), or containing other product of the same category (the fifth column), while CG is able to generate indoor backgrounds consistently. This indicates that CG is able to summarize the rule that the refrigerator is better placed in the room in the advertising image.
5.3.2 Personalized Background Generation
In this subsection, we consider the fine-grained personalized background generation for some specific brands conditioned on the reference advertising image. Specifically, for each pair of product and reference image from BG-pair, we compare our generated backgrounds with those from the following methods: 1) Blended Diffusion [1]: we leverage the loss between the background features of reference and generated image as the classifier-guidance loss. 2) Paint-by-Example [26]: we encode the reference image as the condition. 3) Stable ${}^{\dagger}$ [16]: we first paste the product on the reference image, and then initialize $z_{T}$ with the noisy latent adjusted reference image. 4) ${}^{\dagger}$ [29]: we initialize $z_{T}$ with the noisy latent reference image.
@@ 5.3.2 Personalized Background Generation
Table 2 reports the comparison (see Ref normal-$\rightarrow$ Self). It could be found that our method surpasses the other state-of-the-art methods by a large margin (4.75 CLIP similarity and 1.23 FID). The superiority of CLIP similarity indicates that our method is able to better mimic the character of the background of another reference image, while the extremely low FID demonstrates that our method could generate new backgrounds consistently following the distribution of the advertising images.
Method Ref $\rightarrow$ Self Self $\rightarrow$ Self
Sim ($\uparrow$) FID ($\downarrow$) Sim ($\uparrow$) FID ($\downarrow$)
Blended Diffusion [1] 63.51 3.62 89.02 2.73
Paint-by-Example [26] 62.82 5.96 81.00 5.99
Stable ${}^{\dagger}$ [2] 57.10 7.92 76.10 7.29
${}^{\dagger}$ [29] 62.78 3.92 87.21 3.27
Ours 68.26 2.39 93.63 1.39
Table 2: Background generation conditioned on the reference image on BG-pair. Sim is short for the CLIP similarity (the generated image v.s. the reference image).
Figure 6: Qualitative Results. More visualizations are in the supplementary materials. Our method could generate backgrounds similar to the reference images in style, layout, elements, etc.
Moreover, we consider a simpler situation, where the product in the reference image is exactly the product requiring the background. Therefore, all the expected elements in the generated background could be spatially aligned with those in the reference image, so that the model does not need to understand the reference image deeply. As shown in Table 2 (see Self normal-$\rightarrow$ Self), our method still achieves the best performance in CLIP similarity and FID. Meanwhile the particularly outstanding metrics inspire us to improve the quality of generated background by selecting reference images containing similar-shaped products.
@@
We further conduct a user study for a more focused and comprehensive comparison. We select five well-performed methods from above experiments, average the rank scores (five stands for the best and one for worst) and show the results in Table 3. It could be found that from the participant’s subjective perspective, almost every time the background generated by our method is picked as the most similar one. Moreover, we are encouraged that our method also ranks first in quality. This is more than challenging as pursuing the similarity in the layout and elements might hurt the harmony, especially when the shapes of the product exhibit great differences in the two images. Besides, even without PG, our method outperforms other counterparts in quality, which also proves the effect of the category commonality.
@@
${}^{\dagger}$ [29] 2.39 2.61
Blended Diffusion [1] 2.53 2.77
Paint-by-Example [26] 2.03 2.35
[…]
Ours 4.88 3.84
[…]
We further visualize the qualitative comparison in Fig. 6. It could be found from the first row that our method could mimic the abstract style of the reference image and generate a similar background adaptively. Meanwhile when the background of the reference image is a concrete scene (the last two rows), our method is able to generate background similar in layout, elements, etc.
@@
Table 4: Ablation study on different modules of our method.
5.4 Ablation Study
Effectiveness of CG: To evaluate the effect of the category commonality on personalized background generation, we replace CG with the original ControlNet inpainting model. Then a significant performance deterioration is observed from line (a) and (d) in Table 4. It shows the necessity of the category commonality even with a reference image given, as it forms the basis of the generated background.
Effectiveness of PG: We remove PG and initialize $z_{T}$ with the latent reference image with noise $\epsilon\sim\mathcal{N}(0,1)$ added for $T$ times, to delve into the role of PG in extracting personalized information. From line (b) and (d) in Table 4, this results in the drop of CLIP similarity by $0.76\%$, and the increase of FID by 0.37. It is because the model fails to recognize and leverage the background information explicitly from the reference image.
Impact of Initial Latent Images: We try not initializing $z_{T}$ during inference to study the effect of initializing $z_{T}$ with the noisy reference image. As shown in line (c) and (d) in Table 4, this has a negative influence on the similarity and FID. Thus we conclude that a proper initialization could benefit the quality of generated images.
Comparison on Different Augmentation Strategies: As illustrated in the 3rd column in Fig. 7, when no perturbation strategy is adopted, an anomaly could be observed in where the product in the reference image is originally placed (see the red rectangle). This is alleviated when we employ mask dilation during training. However, as shown in the 4th column in Fig. 7, some elements still look like blunt copies (see the red rectangle). When the mixup strategy is further added, the generated background is finally similar but not identical to the reference background.
@@ 5.4 Ablation Study
6 Conclusion
In this work, we focus on several practical challenges in the e-commerce product background generation. Firstly we improve the efficiency of large-scale generation by a Category-Wise Generator, where we map the common style of each category into a unique identifier by the mask-guided cross attention layers. Furthermore we maintain the personalized styles for some specific brands from reference images effectively. We propose a Personality-Wise Generator and adopt a background perturbation strategy to prevent the copy-and-paste. Finally, we release the first large-scale product background generation dataset BG60k. Experiments show that our method could generate high-quality backgrounds for products from different categories, and backgrounds similar in personalized style to reference images if given.
Appendix A Detailed Analysis on BG60k
The distribution of the top 100 categories in BG60K is illustrated in Fig. 8. It could be found that the categories in BG60K follow the long-tail challenging distribution. It results from the strategy in which we construct BG60k. When collecting images for BG60k, we sort the products in descending order of impressions within one week for better image quality, and retain the top products to ensure the total impressions over 500 million. The high CTR advertising images are then obtained only from these products. Therefore, the final distribution of categories in BG60k follows the real long-tailed category distribution of images with high impressions on the E-commerce platform.
Appendix B Diverse Qualitative Results
In this section we show more diverse qualitative results. Firstly, given the same product, we visualize the generated images conditioned on the different reference images. As illustrated in Fig. 9, all of the generated images share the similar style to the corresponding reference images, which demonstrates the flexibility of our method.
Meanwhile, given the same reference image, we visualize the images generated for different products. As shown in Fig. 10, the generated backgrounds are consistent in layout and elements, but have slight difference in some details. It turns out that our method is able to generated proper backgrounds for different products adaptively, without lost of the personalized styles of the reference backgrounds.
Finally, we provide more qualitative comparisons on background generation conditioned on the reference image in Fig. 11 (see the next page). Our method shows great superiority in generating both artistically stylized backgrounds (the first two rows) and photorealistic backgrounds (the last four rows).
Figure 8: The distribution of the top 100 categories in BG60k. Due to the limited space, we denote each category as an temporary id. The mapping between the category name and id will be released with BG60k.
Figure 9: (a) Product image. (b) Different generated images conditioned on different reference images.
Figure 10: (a) Reference image. (b) Different generated images for different products.
Figure 11: More qualitative comparisons.

### v2

@@ III BG60k
Figure 3: (a) Some unsatisfying examples in LAION-5B [29]. (b) Examples in BG60k, where the training images are more appealing and exquisite.
It is commonly known that the performance of the generation models relies heavily on the quality of the training data. The most widely-used publicly available dataset, like LAION-5B [29] and its subsets, are not designed specifically for e-commerce scenarios. Hence several issues arise in the training images, such as non-salient foreground, the improper perspective, or the dirty and dusky background, as shown in Fig. 3 (a). Models trained on such dirty data will fail to generate satisfying advertising images, as shown in Fig. 4 (b).
To reduce this gap, we release the first large-scale e-commerce background generation dataset BG60k. We first mine millions of advertising images from the famous e-commerce platform JD.COM, then clean them by removing images containing texts or with low CTR. This ends up with 63293 advertising images from 2032 categories. Each item in the dataset is composed of an original advertising image $I$, a mask of the product $M$ and the category of the product. Examples of BG60K are illustrated in Fig. 3 (b). Moreover, we choose 1k products and their original backgrounds, namely BG1k, to compare the difference between the generated and original background statistically. We also select 1.6k pairs of product and reference image to evaluate the ability of generating with personalized style, namely BG-pair.
IV Experiments
IV-A Implementation Details and Evaluation Metrics
Our model is trained on BG60k for 300 epochs. The batch size is 128 and each image is resized to $512\times 512$,. The $\sigma$ in mixup is sampled from a uniform distribution $\mathcal{U}(0.75,0.95)$.
Similar to prior works, we adopt the FID score [30] and the CLIP similarity between generated and original advertising images to measure the quality of the generated images. Besides, a user study is conducted for a more comprehensive comparison, where fifty experienced advertising practitioners are invited to rank the overall aesthetics of generated images and the background consistency.
IV-B Comparison with State-of-the-Art
@@
Large-Scale Background Generation We remove PG and only evaluate CG as no reference images included. As summarized in Table I, CG has the superiority of 1.25 CLIP similarity and 1.85 FID. Thus the backgrounds generated by CG are more like the original high-quality backgrounds and thus better meet the category-wise requirements of the advertisements.
Furthermore, we visualize the features of backgrounds from CG and ControlNet by tSNE in Fig. 4 (a). It could be seen that the features from CG are clustered around the corresponding centers more compactly. Then taking ’refrigerator’ as an example, we compare the generated images in Fig. 4 (b). It turns out that ControlNet might generate backgrounds lacking photorealism or containing other product of the same category, while CG is able to generate indoor backgrounds consistently.
Figure 4: (a) tSNE visualization, where the circle represents the cluster center. Each triangle/square indicates the embedding of one image generated by CG/ControlNet. (b) Visualization of comparison on the category refrigerator.
TABLE I: Comparison conditioned on the prompt template
@@
Personalized Background Generation For each pair of product and reference advertising image from BG-pair, we compare our generated backgrounds conditioned on the reference image with those from the state-of-the-art image-based inpainters [18, 28, 17, 20] in Table II111 and mean initializing $z_{T}$ with the noisy latent reference image. The significant superiority of CLIP similarity indicates that our method could better mimic the character of the reference image, while the extremely low FID demonstrates that our method generates backgrounds consistently following the distribution of the advertising images.
@@
We further conduct a user study for a more focused and comprehensive comparison. We average the rank scores (five stands for the best and one for worst) and show the results in Table III. From the participant’s subjective perspective, almost every time the background generated by our method is picked as the most similar one. Moreover, our method also ranks first in quality. This is more than challenging as pursuing the similarity in the layout and elements might hurt the harmony.
TABLE II: Background generation conditioned on the reference image
Method Ref $\rightarrow$ Self
CLIP Similarity ($\uparrow$) FID ($\downarrow$)
Blended Diffusion [18] 63.51 3.62
Paint-by-Example [28] 62.82 5.96
Stable Diffusion [19] 57.10 7.92
ControlNet [20] 62.78 3.92
Ours 68.26 2.39
TABLE III: User study on background similarity and image quality.
@@
ControlNet [20] 2.39 2.65
Blended Diffusion [18] 2.49 2.71
Paint-by-Example [28] 2.05 2.23
Ours w/o PG 3.16 3.49
Ours 4.88 3.90
TABLE IV: Ablation study on different modules of our method.
@@
IV-C Ablation Study
Effectiveness of CG: We replace CG with the original ControlNet inpainting model and then a significant performance deterioration is observed from line (a) and (d) in Table IV. It shows the necessity of the category commonality even with a reference image given, as it forms the basis of the background.
Effectiveness of PG: We remove PG and initialize $z_{T}$ with the latent reference image with noise $\epsilon\sim\mathcal{N}(0,1)$ added for $T$ times. From line (b) and (d) in Table IV, this results in the drop of CLIP similarity by $0.76\%$, and the increase of FID by 0.37. It is because the model fails to recognize and leverage the background information explicitly from the reference image.
Impact of Initial Latent Images: We try not initializing $z_{T}$ with the noisy latent image, then as shown in line (c) and (d) in Table IV, it has a negative influence on both metrics, proving the importance of the proper initialization o image quality.
Impact of Augmentation Strategies: As illustrated in the 3rd column in Fig. 5, when no perturbation adopted, an anomaly could be observed in where the product is originally placed. With the dilation added, some elements still look like blunt copies. With the mixup strategy added, the generated background is similar but not identical to the reference background.
@@ IV-C Ablation Study
Qualitative Comparison: As shown in Fig. 6, our method could mimic both the abstract style (1st row) and concrete scene (the last two rows) of the reference image better.
Figure 6: Qualitative Comparison.
V Conclusion
In this work, we improve the efficiency of large-scale generation by a Category-Wise Generator, and maintain the personalized styles for some specific brands from reference images effectively by a Personality-Wise Generator. We also release the first large-scale background generation dataset. Experiments show the comprehensive superiority of our methods.